# Complexity reduction of physical models: an equation-free approach by means of scaling

### v1

@@ Appendix B Theorems
Assuming Property A.3, there exists at least a vector $\vec{t}\coloneqq\left(t_{1,1},\dots,t_{N_{\mathsf{p}},N_{\mathsf{x}}}\right)^{T}\in\mathbb{R}^{N_{\mathsf{p}}\,N_{\mathsf{x}}}$ of exponents that ensures (A.3). Under Properties A.1, A.2, A.3, A.4, A.5 and A.6, Theorem B.5 guarantees that such a vector $\vec{t}$ (B.24) satisfies the linear system $\textbf{Y}\,\vec{t}=\vec{\gamma}$ (B.25). Then, there must be at least a solution to such a matrix equation (B.25). Equivalently, the vector $\vec{\gamma}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{p}}}$ must belong to the column space of the matrix $\textbf{Y}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})}$, which is the orthogonal complement of the null space (kernel) of $\textbf{Y}^{T}$, i.e.,
$\vec{\gamma}\in\mathrm{colsp}(\textbf{Y})\equiv\left(\mathrm{ker}(\textbf{Y}^{T})\right)^{\bot}.$ (B.41)
@@
$\textbf{W}\coloneqq\begin{pmatrix}\vec{w}_{1}&\vec{w}_{2}&\cdots&\vec{w}_{N_{\mathsf{r}}}\end{pmatrix}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times N_{\mathsf{r}}}.$ (B.42)
For the matrix W (B.42), the statement (B.41) is equivalent to
$\textbf{W}^{T}\,\vec{\gamma}=\vec{0},$ (B.43)
@@
or, given $\vec{\gamma}$ as in (B.38) and $w_{i,j}$ as the entry of row $i$ and column $j$ of $\textbf{W}^{T}\in\mathbb{R}^{N_{\mathsf{r}}\times(N_{\mathsf{l}}\,N_{\mathsf{p}})}$, to
@@
$\begin{pmatrix}w_{1,1}&w_{1,2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{p}}}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{p}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},1}&w_{N_{\mathsf{r}},2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{p}}}\end{pmatrix}\begin{pmatrix}\vec{\alpha}-\vec{\omega}\\ \vec{\tau}\end{pmatrix}=\vec{0},$ (B.44)
@@
where $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (B.39) is the vector of exponents in (B.20), while $\vec{\omega}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ and $\vec{\tau}\in\mathbb{R}^{N_{\mathsf{l}}\,(N_{\mathsf{p}}-N_{\mathsf{0}})}$ are explicitly known, as specified by (B.37) and (B.33), respectively. By using $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})$ from Theorem B.3 under Properties A.4, A.5 and A.6, we remark that $N_{\mathsf{c}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{0}}=N_{\mathsf{l}}\,(N_{\mathsf{p}}-\mathrm{rank}(\textbf{M}))<N_{\mathsf{l}}\,N_{\mathsf{p}}$, being M (B.8) defined by Theorem B.2 under Property A.6. The latter inequality follows from Property A.6 guaranteeing that M (B.8) cannot be a matrix whose entries are all equal to zero, i.e., $\mathrm{rank}(\textbf{M})\geq 1$. Then, it is possible to write (B.44) in terms of $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$ as the matrix equation $\textbf{S}\,\vec{\alpha}=\vec{s}$ (B.40), where the matrix S and the vector $\vec{s}$ are
@@
$\textbf{S}=\begin{pmatrix}w_{1,1}&w_{1,2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{0}}}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{0}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},1}&w_{N_{\mathsf{r}},2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{r}}\times(N_{\mathsf{l}}\,N_{\mathsf{0}})},$ (B.45)
@@
$\vec{s}=\begin{pmatrix}w_{1,1}&w_{1,2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{0}}}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{0}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},1}&w_{N_{\mathsf{r}},2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}}\end{pmatrix}\vec{\omega}-\begin{pmatrix}w_{1,N_{\mathsf{l}}N_{\mathsf{0}}+1}&w_{1,N_{\mathsf{l}}N_{\mathsf{0}}+2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{p}}}\\ w_{2,N_{\mathsf{l}}N_{\mathsf{0}}+1}&w_{2,N_{\mathsf{l}}N_{\mathsf{0}}+2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{p}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}+1}&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}+2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{p}}}\end{pmatrix}\vec{\tau}\in\mathbb{R}^{N_{\mathsf{r}}},$ (B.46)
@@
with $w_{i,j}$ being the explicitly known entry of row $i$ and column $j$ of the transpose of the matrix W (B.42), while $\vec{\omega}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ and $\vec{\tau}\in\mathbb{R}^{N_{\mathsf{l}}\,(N_{\mathsf{p}}-N_{\mathsf{0}})}$ being explicitly given by (B.37) and (B.33), respectively.
Theorem B.7 enables the computation of numerical values of the exponents $\alpha_{i,k}$ in (B.20) that minimize the deviation $C_{U}$ from unity of the magnitudes of $\lambda_{i}$, with $i\in U\subseteq\{1,\dots,N_{\mathsf{l}}\}$.
Theorem B.7.
Given any fixed choice of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$ and set $U\subseteq\{1,\dots,N_{\mathsf{l}}\}$, the values of $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$ (B.39), $N_{\mathsf{c}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{0}}$, attaining the global minimum of the function
$C_{U}(\vec{\alpha})\coloneqq\sum_{i\in U}\left(\log_{10}(\lambda_{i})\right)^{2},\quad\lambda_{i}=\pi_{1}^{\alpha_{i,1}}\,\pi_{2}^{\alpha_{i,2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{i,N_{\mathsf{0}}}},\quad C_{U}:\mathbb{R}^{N_{\mathsf{c}}}\to[0,\infty),$ (B.47)
subject to a constraint of the form
$\boldsymbol{\Sigma}\,\vec{\alpha}=\vec{\sigma},\quad\boldsymbol{\Sigma}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}\times N_{\mathsf{c}}},\quad\vec{\sigma}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}},\quad\hat{N_{\mathsf{r}}}\in\mathbb{N},$ (B.48)
satisfy the system of equations
$\boldsymbol{\Sigma}\,\vec{\alpha}=\vec{\sigma},\quad\nabla C_{U}(\vec{\alpha})\,+\,\boldsymbol{\Sigma}^{T}\,\vec{\nu}=\vec{0},$ (B.49)
assuming that
$\mathrm{rank}(\boldsymbol{\Sigma})=\hat{N_{\mathsf{r}}}<N_{\mathsf{c}},$ (B.50)
and there is a vector $\vec{\nu}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}}$ such that (B.49) holds.
@@
The twice continuously differentiable function $C_{U}=C_{U}(\vec{\alpha})$ in (B.47) is convex as its Hessian matrix is diagonal with non-negative entries $\forall\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$. Suppose one also aims to satisfy a constraint of the form (B.48). It follows that finding any minimum of $C_{U}=C_{U}(\vec{\alpha})$ reads as a convex optimization problem with equality constraints, as considered in Section 10.1 of [20].
Then, as guaranteed in Section 10.1 of [20] by assuming (B.50), the vector $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$ is optimal for the problem under consideration if and only if there is a vector $\vec{\nu}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}}$ such that (B.49) holds. Any found local minimizer is automatically a global minimizer, because the considered problem is convex, as discussed in Section 4.2.2 of [20].

### v2

@@ Appendix E Proof of Theorem 5
Bearing Assumption˜3, there exists at least a vector $\vec{t}\coloneqq\left(t_{1,1},\dots,t_{N_{\mathsf{p}},N_{\mathsf{x}}}\right)^{T}\in\mathbb{R}^{N_{\mathsf{p}}\,N_{\mathsf{x}}}$ of exponents that ensures (16). Under Assumptions˜1, 2, 3, 4, 5 and 6, Theorem 4 guarantees that such a vector $\vec{t}$ (23) satisfies the linear system $\textbf{Y}\,\vec{t}=\vec{\gamma}$ (24). Then, there must be at least a solution to such a matrix equation (24). Equivalently, the vector $\vec{\gamma}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{p}}}$ must belong to the column space of the matrix $\textbf{Y}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})}$, which is the orthogonal complement of the null space (kernel) of $\textbf{Y}^{T}$, i.e.,
$\vec{\gamma}\in\mathrm{colsp}(\textbf{Y})\equiv\left(\mathrm{ker}(\textbf{Y}^{T})\right)^{\bot}.$ (E.1)
@@
$\textbf{W}\coloneqq\begin{pmatrix}\vec{w}_{1}&\vec{w}_{2}&\cdots&\vec{w}_{N_{\mathsf{r}}}\end{pmatrix}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times N_{\mathsf{r}}}.$ (E.2)
For the matrix W (E.2), the statement (E.1) is equivalent to
$\textbf{W}^{T}\,\vec{\gamma}=\vec{0},$ (E.3)
@@
or, given $\vec{\gamma}$ as in (D.13) and $w_{i,j}$ as the entry of row $i$ and column $j$ of $\textbf{W}^{T}\in\mathbb{R}^{N_{\mathsf{r}}\times(N_{\mathsf{l}}\,N_{\mathsf{p}})}$, to
@@
$\begin{pmatrix}w_{1,1}&w_{1,2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{p}}}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{p}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},1}&w_{N_{\mathsf{r}},2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{p}}}\end{pmatrix}\begin{pmatrix}\vec{\alpha}-\vec{\omega}\\ \vec{\tau}\end{pmatrix}=\vec{0},$ (E.4)
@@
where $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (25) is the vector of exponents in (22), while $\vec{\omega}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ and $\vec{\tau}\in\mathbb{R}^{N_{\mathsf{l}}\,(N_{\mathsf{p}}-N_{\mathsf{0}})}$ are explicitly known, as specified by (D.12) and (D.8), respectively. By using $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})$ from Theorem 2 under Assumptions˜4, 5 and 6, we remark that $N_{\mathsf{c}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{0}}=N_{\mathsf{l}}\,(N_{\mathsf{p}}-\mathrm{rank}(\textbf{M}))<N_{\mathsf{l}}\,N_{\mathsf{p}}$, being M (20) defined by Theorem 1 under Assumption˜6. The latter inequality follows from Assumption˜6 guaranteeing that M (20) cannot be a matrix whose entries are all equal to zero, i.e., $\mathrm{rank}(\textbf{M})\geq 1$. Then, it is possible to write (E.4) in terms of $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$ as the matrix equation $\textbf{S}\,\vec{\alpha}=\vec{s}$ (26), where the matrix S and the vector $\vec{s}$ are
@@
$\textbf{S}=\begin{pmatrix}w_{1,1}&w_{1,2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{0}}}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{0}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},1}&w_{N_{\mathsf{r}},2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{r}}\times(N_{\mathsf{l}}\,N_{\mathsf{0}})},$ (E.5)
@@
$\vec{s}=\begin{pmatrix}w_{1,1}&w_{1,2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{0}}}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{0}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},1}&w_{N_{\mathsf{r}},2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}}\end{pmatrix}\vec{\omega}-\begin{pmatrix}w_{1,N_{\mathsf{l}}N_{\mathsf{0}}+1}&w_{1,N_{\mathsf{l}}N_{\mathsf{0}}+2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{p}}}\\ w_{2,N_{\mathsf{l}}N_{\mathsf{0}}+1}&w_{2,N_{\mathsf{l}}N_{\mathsf{0}}+2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{p}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}+1}&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}+2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{p}}}\end{pmatrix}\vec{\tau}\in\mathbb{R}^{N_{\mathsf{r}}},$ (E.6)
@@
with $w_{i,j}$ being the explicitly known entry of row $i$ and column $j$ of the transpose of the matrix W (E.2), while $\vec{\omega}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ and $\vec{\tau}\in\mathbb{R}^{N_{\mathsf{l}}\,(N_{\mathsf{p}}-N_{\mathsf{0}})}$ being explicitly given by (D.12) and (D.8), respectively.
Appendix F Proof of Theorem 6
See 6
@@ Appendix F Proof of Theorem 6
By employing the notation (1) in (14) (Assumption˜1), one has
$\lambda_{i}=\tilde{\kappa}_{i}\,\tilde{\theta}_{1}^{c_{i,1}}\,\tilde{\theta}_{2}^{c_{i,2}}\,\cdots\,\tilde{\theta}_{N_{\mathsf{x}}}^{c_{i,N_{\mathsf{x}}}}=\kappa_{i}[\tilde{\kappa}_{i}]\,\theta_{1}^{c_{i,1}}[\tilde{\theta}_{1}]^{c_{i,1}}\,\theta_{2}^{c_{i,2}}[\tilde{\theta}_{2}]^{c_{i,2}}\,\cdots\,\theta_{N_{\mathsf{x}}}^{c_{i,{N_{\mathsf{x}}}}}[\tilde{\theta}_{N_{\mathsf{x}}}]^{c_{i,{N_{\mathsf{x}}}}},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (F.1)
$\lambda_{i}]=[\tilde{\kappa}_{i}]\,[\tilde{\theta}_{1}]^{c_{i,1}}\,[\tilde{\theta}_{2}]^{c_{i,2}}\,\cdots\,[\tilde{\theta}_{N_{\mathsf{x}}}]^{c_{i,{N_{\mathsf{x}}}}}=1,\quad\forall i=1,\dots,N_{\mathsf{l}},$ (F.2)
it follows from (14) and (F.1) that
$\lambda_{i}=\kappa_{i}\,\theta_{1}^{c_{i,1}}\,\theta_{2}^{c_{i,2}}\,\cdots\,\theta_{N_{\mathsf{x}}}^{c_{i,{N_{\mathsf{x}}}}}\in(0,\infty),\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (F.3)
Then, (F.3) allows writing the function $C_{U}$ as
$C_{U}(\vec{\rho})=\sum_{i\in U}\left(\log_{10}(\kappa_{i})+\sum_{k=1}^{N_{\mathsf{x}}}\,c_{i,k}\,\rho_{k}\right)^{2},\quad\forall\vec{\rho}\coloneqq(\rho_{1},\dots,\rho_{N_{\mathsf{x}}})\in\mathbb{R}^{N_{\mathsf{x}}},$ (F.4)
where $\rho_{k}\coloneqq\log_{10}(\theta_{k})$ for any $k=1,\dots,N_{\mathsf{x}}$. The twice-differentiable function $C_{U}(\vec{\rho})$ (F.4) is convex because its Hessian matrix is positive semi-definite $\forall\vec{\rho}\in\mathbb{R}^{N_{\mathsf{x}}}$. Then, as guaranteed by Theorem 2.5 in [Nocedal_book_2006], the global minimizers of $C_{U}$ can be found by imposing $\partial_{\rho_{j}}C_{U}=0$, $\forall j=1,\dots,N_{\mathsf{x}}$, yielding to the linear system (28) of $N_{\mathsf{x}}$ equations with $N_{\mathsf{x}}$ unknowns $\rho_{1},\dots,\rho_{N_{\mathsf{x}}}\in\mathbb{R}$.
Appendix G Proof of Theorem 7
See 7
Proof.
The twice continuously differentiable function $C_{U}=C_{U}(\vec{\alpha})$ in (29) is convex as its Hessian matrix is diagonal with non-negative entries $\forall\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$. Suppose one also aims to satisfy a constraint of the form (30). It follows that finding any minimum of $C_{U}=C_{U}(\vec{\alpha})$ reads as a convex optimization problem with equality constraints, as considered in Section 10.1 of [boyd_vandenberghe_2004].
Then, as guaranteed in Section 10.1 of [boyd_vandenberghe_2004] by assuming (32), the vector $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$ is optimal for the problem under consideration if and only if there is a vector $\vec{\nu}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}}$ such that (31) holds. Any found local minimizer is automatically a global minimizer, because the considered problem is convex, as discussed in Section 4.2.2 of [boyd_vandenberghe_2004].